What is the equation of the line with y-intercept 7 and x-intercept 5?
Answer by fractalier(6550)

The "intercept" form of a line is not often well-taught...it is
x/a + y/b = 1
where a and b are the x- and y- intercepts, respectively.
Thus we get
x/5 + y/7 = 1
To put this in standard form, we multiply by 35 and get
7x + 5y = 35